# VR-Robo: A Real-to-Sim-to-Real Framework for Visual Robot Navigation and Locomotion

### v1

@@ IV-C RL in Reconstructed Simulation
State Space: The high-level policy actor network receives four types of inputs: 1) RGB Image Feature ($I_{a}$): We employ a frozen, pre-trained Vision Transformer (ViT) [12] to extract features from RGB images. ViT’s attention mechanism is well-suited for detecting positions and specific colors. This approach compresses the input size significantly and accelerates training. 2) RGB Command ($C_{a}$): A RGB vector indicating the desired color of the target cone. For example, $[1,0,0]$ represents the red cone. 3) Last Action: The action output by the policy during the previous timestep. 4) Proprioception ($P_{a}$): This includes the robot’s base angular velocity, projected gravity, joint positions, and joint velocities. We use an asymmetric actor-critic structure. The critic’s observation includes the actor’s observation (with noise removed) as well as the robot’s world coordinate position and orientation, and the target cone’s world coordinate position. This design improves the estimation of value function and enhances the training process.
@@
$r_{\text{reach\_goal}}=\begin{cases}R_{\text{max}},&\text{if }\|\mathbf{p}_{\text{robot}}-\mathbf{p}_{\text{goal}}\|_{2}\leq\epsilon,\\ 0,&\text{otherwise}.\end{cases}$ (14)
@@
$d$ $=\|\mathbf{p}_{\text{robot}}-\mathbf{p}_{\text{goal}}\|_{2}.$ (16)
@@
$d_{z}$ $=|z_{\text{robot}}-z_{\text{goal}}|.$ (18)
@@
$r_{\text{goal\_heading}}=-|\psi_{\text{robot}}-\psi_{\text{goal}}|,$ (19)
@@
In addition, regularization rewards are incorporated to further refine the robot’s performance and ensure stable and efficient behavior. It contains Stop_at_goal, Track_lin_vel, Track_ang_vel, and Action_l2. The detailed mathematical expressions and weights of all rewards can be found in Appendix -D.
@@ V Experiments
To evaluate the effectiveness of VR-Robo, we construct two Real-to-Sim environments by reconstructing 3D scenes from phone-captured image collections. These two environments capture an indoor scene with complex terrains. Our experiments are designed to evaluate two main aspects of VR-Robo: 1) its ability to reconstruct realistic and interactive simulation environments with GS-mesh hybrid representation, supporting the locomotion policy training through both visual observations and mesh interactions; 2) its capability to minimize the Sim-to-Real gap when deployed to the real world. We conducted extensive experiments in both simulation and real-world settings to validate these capabilities.
@@ V-A Experiment Setup
Real-to-Sim Reconstruction. We reconstruct two distinct indoor room environments, each characterized by specific terrains. We also randomize the environments by incorporating three cones colored red, green, and blue. We use iPads or iPhones to take photos, which are easily available. Detailed procedures for image acquisition and coordinate alignment are provided in the Appendix -E.
Simulated Locomotion Training. We conduct policy training in Isaac Sim using a single 48GB NVIDIA RTX 4090D GPU. For the low-level policy, we instantiate 4,096 quadruped robot agents in parallel and train for 80,000 iterations from scratch. The low-level policy operates at a frequency of 50 Hz. In contrast, the high-level policy is trained from scratch over 8,000 iterations with 64 quadruped agents. The high-level policy operates at a frequency of 5 Hz. The entire training process requires approximately three days to complete. For visual encoding, we employ the “vit_tiny_patch16_224” Vision Transformer model, omitting its final classification head. The Isaac Sim simulator and the 3DGS rasterization-based renderer communicate via a TCP network.
Sim-to-Real Deployment. We deploy our policy on the Unitree Go2 quadruped robot, which is equipped with an NVIDIA Jetson Orin Nano (40 TOPS) as the onboard computer. We use ROS2 [30] for communication between the high-level policy, low-level policy, and the robot. Both policies run onboard, with the high-level policy operating at 5 Hz and the low-level policy running at 50 Hz. The robot receives desired joint positions from the low-level policy for PD control ($K_{p}=40.0$, $K_{d}=1.0$). We use an Insta360 Ace camera to capture RGB images. The camera captures images at a resolution of $320\times 180$. After calibration, it has a horizontal field of view (FOVX) of 1.5701 radians and a vertical field of view (FOVY) of 1.0260 radians.
V-B Simulation Experiments
  • Depth Policy: Following a previous parkour approach [61], we replace the RGB input with a depth image and adopt a simple CNN as the vision encoder.
For ablation experiments,
  • Texture Mesh: We incorporate a SuGaR [16]-reconstructed textured mesh into Isaac Sim and render RGB observations directly from this mesh.
  • CNN Encoder: We replace the ViT vision encoder with a CNN image encoder, specifically “MobilenetV3” [18], and remove the last classification layer.
Figure 7: Drawbacks of the Texture Mesh and Depth Input Images. (Left: agent ego-centric observation; Right: simulation environment in bird-eye-view.) The mesh is cropped for use in the simulator due to the large original size. As shown in the area marked by the red box, the mesh exhibits bulges in texture-less regions, such as the ground and walls, which hinder the robot’s movement. Additionally, controlling the lighting in the simulator poses a significant challenge. The depth map is rendered directly from the simulator but cannot distinguish the color of the cones.
The visualize explanations of Texture Mesh and Depth are shown in Figure 7. We evaluated all methods in a ‘red cone reaching’ task within Scene 1. The red cone is randomly placed in one of three regions: left (ground), middle (terrain), or right (ground). We generate 150 distinct cone placements, specified prior to any experiments, and apply the same setup for all methods to ensure a fair comparison.
TABLE I: Comparison results in the simulation setting.
Method SR $\uparrow$ ART (s) $\downarrow$
Ours 100.00% 4.94
Imitation Learning (IL) 8.67% 14.01
Depth Policy 6.00% 14.13
TABLE II: Ablation results in the simulation setting.
Method SR $\uparrow$ ART (s) $\downarrow$
Ours 100.00% 4.94
Texture Mesh 22.00% 12.73
CNN Encoder 54.67% 10.04
TABLE III: Comparison and ablation experimental results in the real-world setting.
@@ V-B Simulation Experiments
The comparison and ablation results are shown in Table I and Table II separately. Our method consistently attains higher performance across all evaluation metrics by a large margin. In particular, the Imitation Learning (IL) baseline suffers from insufficient data samples and lacks policy exploration. The depth policy fails to distinguish between different cones and cannot be effectively trained, resulting in behavior akin to random actions. Similarly, the texture mesh is unable to generate realistic images, and the CNN encoder struggles to extract precise features from the images. These findings collectively confirm the effectiveness and robustness of our proposed approach in diverse experimental setups.
@@ V-C Real-World Experiments
In addition to the methods mentioned above, we conduct two additional baseline methods for real-world experiments. These methods are specifically designed in the real-world setting and cannot be directly compared in simulation:
@@
[…]
Figure 8: Tasks of different difficulties (easy, medium, and hard) and different positions (left, middle, and right).
The quantitative results are presented in Table III. Our method achieves the highest success rates across all difficulty levels and consistently records the shortest average reaching times, demonstrating its efficiency and robustness. Among the other methods, SARO achieves moderate success on easy tasks (66.67%) but performs poorly on medium and hard tasks. This underperformance is primarily attributed to the lack of historical context and limited exploration capability in complex scenarios. If the robot cannot initially see the goal, it is unlikely to succeed. Furthermore, the absence of domain randomization significantly reduces SARO’s effectiveness, primarily due to the large sim-to-real gap with RGB-only observations.
We further conduct qualitative experiments under varying conditions, including various target cone colors, changes in lighting conditions, different scenes, random interference, background variations, and training on the Unitree G1 humanoid robot, as shown in Figure 9. A detailed demonstration is provided in the supplementary materials. These experiments highlight the robustness of our method, showcasing its adaptability to a wide range of environments and conditions.
Figure 9: Qualitative experiments under varying conditions.
@@ VI-A Limitations
1) Simple Scene and Task: while VR-Robo facilitates easier real-to-sim transfer through its GS-mesh hybrid representation and coordinate alignment workflow, our current simulation environment is restricted to indoor scenes with static terrains. Future work could explore the simulation of large-scale and complex scenarios, such as factory settings or outdoor street environments, incorporating both static and dynamic obstacles to support a wider range of downstream tasks. 2) Workflow Efficiency: despite the accelerated visual locomotion policy training afforded by the real-time rendering capabilities of 3DGS, training a large number of agents in parallel remains relatively slow due to the serial rendering pipeline across agents and the significant communication costs between the simulator and the GS renderer. This also limits training on finding cones of different colors at the same time. Future efforts may focus on designing CUDA-based parallel randomization and rendering pipelines, as well as more efficient interaction methods between the GS and the simulator. 3) Model Pretraining: as discussed in  subsection V-A, VR-Robo currently requires approximately three days to train a reinforcement learning (RL) locomotion policy from scratch for each task. We anticipate that more effective pretraining methods could mitigate this time bottleneck.
@@ VI-B Conclusion
We introduce VR-Robo, a real-to-sim-to-real framework designed to train visual locomotion policies within a photorealistic and physically interactive simulation environment, thereby minimizing the sim-to-real gap. Extensive experimental results demonstrate that by integrating real-world environments into the simulator with randomization, agents successfully learn robust and effective policies for challenging high-level tasks and can be zero-shot deployed to real-world scenarios. In future work, we aim to extend our framework to larger-scale and more complex environments, incorporating diverse agents such as humanoid robots and mobile manipulators.

### v3

@@ IV-C Reinforcement Learning in Reconstructed Simulation
State Space: The high-level policy actor network receives four types of inputs: 1) RGB Image Feature ($I_{a}$): We employ a frozen, pre-trained Vision Transformer (ViT)[38] to extract features from RGB images. ViT’s attention mechanism is well-suited for detecting positions and specific colors. This approach compresses the input size significantly and accelerates training. 2) RGB Command ($C_{a}$): A RGB vector indicating the desired color of the target cone. For example, $[1,0,0]$ represents the red cone. 3) Last Action: The action output by the policy during the previous timestep. 4) Proprioception ($P_{a}$): This includes the robot’s base angular velocity, projected gravity, joint positions, and joint velocities. We use an asymmetric actor-critic structure. The critic’s observation includes the actor’s observation (with noise removed) as well as the robot’s world coordinate position and orientation, and the target cone’s world coordinate position. This design improves the estimation of the value function and enhances the training process.
@@
$r_{\text{reach\_goal}}=\begin{cases}R_{\text{max}},&\text{if }\|\mathbf{p}_{\text{robot}}-\mathbf{p}_{\text{goal}}\|_{2}\leq\epsilon,\\ 0,&\text{otherwise}.\end{cases}\vspace{-1mm}$ (14)
@@
$d$ $=\|\mathbf{p}_{\text{robot}}-\mathbf{p}_{\text{goal}}\|_{2}.\vspace{-3mm}$ (16)
@@
$d_{z}$ $=|z_{\text{robot}}-z_{\text{goal}}|.\vspace{-3mm}$ (18)
@@
$r_{\text{goal\_heading}}=-|\psi_{\text{robot}}-\psi_{\text{goal}}|,\vspace{-1mm}$ (19)
@@
In addition, regularization rewards are incorporated to further refine the robot’s performance and ensure stable and efficient behavior. It contains Stop_at_goal, Track_lin_vel, Track_ang_vel, and Action_l2.
@@ V Experiments
Our experiments are designed to evaluate two main aspects of VR-Robo: 1) its ability to reconstruct realistic and interactive simulation environments with GS-mesh hybrid representation, supporting the locomotion policy training through both visual observations and mesh interactions; 2) its capability to minimize the Sim-to-Real gap when deployed to the real world. We conducted extensive experiments in both simulation and real-world settings to validate these capabilities.
@@ V-A Experiment Setup
Real-to-Sim Reconstruction. We reconstruct 6 distinct indoor room environments, each characterized by specific terrains. We also randomize the environments by incorporating three cones colored red, green, and blue. We use iPads or iPhones to take photos, which are easily available.
Simulated Locomotion Training. We conduct policy training in Isaac Sim using a single NVIDIA 4090D GPU. For the low-level policy, we instantiate 4,096 quadruped robot agents in parallel and train for 80,000 iterations from scratch. In contrast, the high-level policy is trained from scratch over 8,000 iterations with 64 quadruped agents. The entire training process requires approximately three days to complete. For visual encoding, we employ the Vision Transformer model “vit_tiny_patch16_224”, omitting its final classification head. The Isaac Sim simulator and the 3DGS rasterization-based renderer communicate via a TCP network.
Sim-to-Real Deployment. We deploy our policy on the Unitree Go2 quadruped robot, which is equipped with an NVIDIA Jetson Orin Nano (40 TOPS) as the onboard computer. We use ROS2[39] for communication between the high-level policy, low-level policy, and the robot. Both policies run onboard. The robot receives desired joint positions from the low-level policy for PD control ($K_{p}=40.0$, $K_{d}=1.0$). We use an Insta360 Ace camera to capture RGB images. The camera captures images at a resolution of $320\times 180$. After calibration, it has a horizontal field of view (FOVX) of 1.5701 radians and a vertical field of view (FOVY) of 1.0260 radians.
TABLE I: Comparison and ablation experimental results in the real-world setting.
@@
V-B Simulation Experiments
Figure 8: Training reward comparison. Pretrained CNN [41] encoder struggles to extract precise semantic features, leading to unstable and worse rewards compared to ViT [38] encoder.
Figure 9: Illustration of textured mesh and random background.
TABLE II: Comparison results in the simulation setting.
Method SR $\uparrow$ ART (s) $\downarrow$
Ours 100.00% 4.94
Imitation Learning (IL) 8.67% 14.01
Random Background 43.33% 11.75
TABLE III: Ablation results in the simulation setting.
Method SR $\uparrow$ ART (s) $\downarrow$
Ours 100.00% 4.94
Textured Mesh 22.00% 12.73
CNN Encoder 54.67% 10.04
TABLE IV: Reconstruction quality comparison.
Method Planar 3DGS Textured Mesh
PSNR $\uparrow$ 29.73 27.73
  • Random Background: LucidSim [10] uses generative models to produce visual observations. Since the training and deployment code has not been open-sourced, we re-implemented it by using random images from ImageNet as the background and maintaining the target object.
For ablation experiments,
  • Textured Mesh: We use SuGaR[42] to reconstruct textured mesh and support direct mesh rendering in Isaac Sim as visual observation.
  • CNN Encoder: We replace ViT with a CNN [41] encoder, and remove the last classification layer.
We evaluated all methods in a ‘red cone reaching’ task within Scene 1. The red cone is randomly placed in one of three regions: left (ground), middle (terrain), or right (ground). We generate 150 distinct cone placements, specified before any experiments, and apply the same setup for all methods to ensure a fair comparison.
The comparison and ablation results are shown in Table II and Table III, respectively. Our method consistently achieves higher performance across all evaluation metrics by a large margin. In particular, the Imitation Learning (IL) baseline suffers from insufficient data samples and a lack of policy exploration. As shown in Figure 8, the CNN encoder struggles to extract precise features from the images. In this work, we reconstruct solely from RGB images. The textured mesh obtained from GS exhibits noticeable bulges in texture-less regions such as the ground and walls, as shown in Figure 9 and Table IV. These artifacts degrade the rendering quality and can even hinder the robot’s movement. The quality and resolution of the mesh itself significantly affect rendering fidelity, whereas GS is less sensitive to such factors. In contrast, our method crops and utilizes only the mesh of the central terrain, and renders observation from GS, effectively avoiding these issues. The Random Background setting (the bottom-right image) loses important features specific to the original scene (the bottom-left image). These scene-specific features are crucial for the robot to effectively explore the environment.
@@ V-C Real-World Experiments
We conduct qualitative experiments under varying conditions, including 6 different scenes, various target cone colors, changes in lighting conditions, random interference, background variations, and training on the different robots, as shown in Figure 2. A detailed demonstration is provided in the supplementary materials. These experiments highlight the robustness of our method, showcasing its adaptability to a wide range of environments and conditions.
Figure 10: Tasks of different difficulties (easy, medium, and hard) and different target positions (left, middle, and right).
We further conduct quantitative experiments, as presented in Table I. In addition to the methods mentioned above, we conduct two additional baseline methods for real-world experiments. These methods are specifically designed in the real-world setting and cannot be directly compared in simulation:
@@
All the real-world quantitative experiments is in the ‘red cone reaching’ task setting in Scene 1. We categorize the task into three levels of difficulty based on the robot’s starting position: easy, medium, and hard. For easy tasks, the robot starts with the target cone directly visible. For medium tasks, the robot needs to turn to a certain degree to see the target cone. For hard tasks, the robot starts very far from the cone and faces a nearly opposite direction to the target cone. We randomly select 3 sets of cone positions, with each set containing one easy, one medium, and one hard task, as shown in Figure 10. For each task, we repeat the experiment 5 times. We then calculate the Success Rate (SR) and Average Reaching Time (ART). For SARO, the maximum time is extended to 60 seconds due to the long inference time required by the Vision-Language Model (VLM). In the real robot experiments, we consider the task successful if the robot makes contact with the target cone.
Our method achieves the highest success rates across all difficulty levels and consistently records the shortest average reaching times, demonstrating its efficiency and robustness. Note that due to inherent randomness in the real-world setting, there is a failure case in the medium task, even though all trials succeed in the hard task. Among the other methods, SARO achieves moderate success on easy tasks (66.67%) but performs poorly on medium and hard tasks. This under-performance is primarily attributed to the lack of historical context and limited exploration capability in complex scenarios. If the robot cannot initially see the goal, it is unlikely to succeed.
@@ VI-A Limitations
The current environment is limited to indoor scenes with static terrains, which constrains the diversity of scenarios. Training a locomotion policy from scratch for each task takes approximately three days, posing a significant time burden. Additionally, meshes reconstructed using only RGB input often contain structural imperfections and require manual post-processing, as they lack the fidelity that could be provided by additional sensing modalities such as depth cameras or LiDAR.
@@ VI-B Conclusion
We introduce VR-Robo, a real-to-sim-to-real framework designed to train visual locomotion policies within a photorealistic and physically interactive simulation environment. Extensive experimental results demonstrate that our agents successfully learn robust and effective policies for challenging high-level tasks and can be zero-shot deployed to diverse real-world scenarios. In future work, we aim to extend our framework to larger-scale and more complex environments, incorporating generative models into scene reconstruction for more generalizable policy learning.